# Performance of multiple filter-cavity schemes for frequency-dependent squeezing in gravitational-wave detectors

### v1

@@ Abstract
Gravitational-wave detectors use state-of-the-art quantum technologies to circumvent vacuum fluctuations via squeezed states of light. Future detectors such as Einstein Telescope may require the use of two filter cavities or a 3-mirror, coupled filter cavity to achieve a complex rotation of the squeezing ellipse in order to reduce the quantum noise over the whole detector bandwidth. In this work, we compare the theoretical feasibility and performances of these two systems and their resilience with respect to different degradation sources (optical losses, mismatching, locking precision). We provide both analytical models and numerical insights. We extend previous analysis on squeezing degradation and find that the coupled cavity scheme provides similar or better performances than the two-cavity option, in terms of resilience with respect to imperfections and optical losses. We propose a possible two-step implementation scheme for Einstein Telescope using a single filter cavity that can be possibly upgraded into a coupled filter cavity.
@@ I Introduction
Over the past decade, LIGO, Virgo and KAGRA have opened the era of gravitational-wave (GW) astronomy, detecting hundreds of signals from merging compact binary objects, including black-hole binaries, neutron-star binaries and black-hole-neutron-star systems [1]. These observations have led to a wide range of scientific insights in fields as diverse as general relativity, astrophysics and cosmology [2].
@@
Quantum noise limits the ultimate sensitivity of GW detectors. Such noise is due to vacuum fluctuations of the electromagnetic field entering the detector through the output port. Quantum noise can be mitigated by generating and injecting squeezed vacuum through this port [3, 4]. Because of the inherent frequency response of the interferometer, and the ponderomotive coupling with the mirrors, such a squeezed vacuum state needs to acquire a frequency dependence in order to optimally reduce quantum noise in the whole detector bandwidth. Precisely, the interferometer rotates the squeezed quadratures in a frequency-dependent manner. Fig. 1 shows the simplified optical scheme of such a quantum-enhanced GW interferometer.
@@
Figure 1: Simplified optical scheme for dual-recycled GW interferometer with Frequency-Dependent Squeezing injected through the output port. Current and future generations of GW detectors use the same simplified scheme, but ET-LF operates in a detuned configuration.
@@
Frequency-Dependent Squeezing (FDS) is achieved by first generating a frequency-independent squeezed vacuum state, then passing it through one or several filter cavities to rotate its quadratures, thus compensating the response of the interferometer [5, 6, 7]. Recently, both Advanced Virgo and Advanced LIGO demonstrated FDS using a single filter cavity on site [8, 9], because the interferometers are operated in a broadband configuration [10]. In particular, the implementation of FDS in Advanced LIGO improved the detection rate up to 65% [11] and allowed operation below the standard quantum limit [12]. While not currently implemented in LIGO or Virgo, note that an alternative scheme – called EPR squeezing – which does not directly require a filter cavity to achieve FDS, has been theoretically proposed [13] and demonstrated in proof-of-principle experiments [14, 15].
@@
At present, “third-generation” (3G) detectors, to be built in new infrastructures, are being studied and aim to increase sensitivity by a factor of $\sim 10$ compared to existing detectors. They will expand the science cases of GW, addressing fundamental open questions such as the nature of gravity and dark energy, the properties of nuclear matter, and the formation of neutron stars and black holes throughout cosmic history. Einstein Telescope, one of such 3G detectors, is composed of two interferometer designs: a “hot” detector, optimized for high-frequency sensitivity, and a “cold” detector (ET-LF), operating at cryogenic temperatures for improved low-frequency sensitivity [16]. This work focuses on the quantum noise reduction (QNR) in ET-LF as a 10-km, single L-shaped (right angle between the arms), Michelson interferometer.
@@
While ET-LF’s sensitivity will be limited by quantum noise, similar to current GW detectors, its mitigation will require two filter cavities (referenced as 2FC in the following) [17]. This is because ET-LF will operate in a detuned (non-broadband) configuration [18, 19]. Among alternatives schemes that do not require two filter cavities, such as quantum teleportation [20], we consider here a three-mirror cavity, otherwise known as a coupled filter cavity (henceforth referenced as CFC), which has been first proposed in [21]. A comparison of 2FC vs. CFC in the lossless case was performed in [22], where it was identified that the equivalence between the two schemes imposed a value for the CFC middle mirror transmissivity apparently smaller than current state of the art techniques in mirror coating.
@@
In this paper, we review the equivalency between 2FC and CFC, and address the issue of the middle mirror transmission through compensating mechanisms (Section II). In particular, we show that, for longer filter cavities than that studied in [22], its value is actually readily achievable. We then perform a complete comparison between 2FC and CFC configuration in terms of squeezing degradation budget, in particular including the effects of optical loss, mismatch and phase noise (Section III). It is shown that, theoretically, CFC may provide better resilience to mode-matching and phase noise than 2FC. Finally, we introduce a tuned configuration of ET-LF (with single filter cavity) as an intermediary step before achieving detuned ET-LF, and study its astrophysical merit (Section IV).
@@ II.1 Scaling law between 2FC and CFC
with $\nu_{c}=\frac{c}{2L_{c}}$ the free spectral range (FSR) of the c-cavity, $\gamma_{i}=\frac{cT_{i}}{4L_{i}}$ and $\Delta\omega_{i}$ the linewidth and detuning of cavity labeled $i$.
@@
Table 1 provides the parameters which will be used in the rest of this paper (taken from [23]). We briefly comment on the chosen value of 30 ppm for round trip loss $\Lambda$. They have been estimated according to the empirical scaling law:
Λ=30 $\Lambda=30\text{ ppm}\times\left(\frac{L}{300\text{ m}}\right)^{0.3}\left(\frac{1064\text{ nm}}{\lambda}\right)^{2}$ (5)
where the values of 30 ppm has been found by simulations in [24] for a 300-m cavity illuminated with a 1064 nm laser. For the scaling low with the length of the cavity we use the result of [25]. The scaling with the laser wavelength is given by the physics of the scattering [26]. This can be considered a conservative estimation, since we expect future advances in the mirror and coating techniques.
@@
$\theta_{\text{HD}}$ Homodyne angle $-0.27$ rad
@@
$\Lambda$ Round Trip Loss 30 ppm (typ.)
$\delta L$ FC RMS length noise 1 pm (typ.)
@@
$L_{\text{FC1,2}}$ Length of each FC $5$ km
@@
$L_{\text{FC1,2}}$ Length of each FC $1$ km
$L_{\text{CFC}}$ Total length of CFC $2$ km
@@
Table 1: Parameters used in this study. For comparison, we show the parameters of 2FC and CFC if the total length is either 2 or 10 km.
@@ II.2 The middle mirror in CFC
The middle mirror is a critical component of the CFC configuration. We first comment on concerns raised by the unusually small value for $T_{a}$. First, note that while its value in the $L_{\text{CFC}}=2$ km case ($T_{a}<0.3$ ppm) is indeed much lower than what is implemented in the current generation of GW detectors, the $L_{\text{CFC}}=10$ km case ($T_{a}=6.75$ ppm) is comparable to the transmission of the end mirror of the currently installed filter cavity in Virgo (see [27] page 350), thus technically already achieved for comparable wavelengths. Therefore, in the rest of this paper, unless explicitly specified otherwise, we will work with $L_{\text{CFC}}=10$ km.
Interestingly, if one considers the a-cavity as a single cavity, – that is, the input mirror is $T_{a}$ – then it may seem that such a cavity is loss-dominated (Λ=30 $\Lambda=30\text{ ppm}\gg T_{a}$), and that the finesse would be of the order of $10^{5}$. However, since the a-cavity is embedded in a coupled cavity system, the correct qualitative picture is to consider that the first two mirrors ($T_{c}$ and $T_{a}$) form a cavity, which is equivalent to an effective mirror with variable transmission $T_{\text{eff}}$ given as a function of the sideband frequency $\Omega$ by
@@
[…]
Small deviations from the optimal middle mirror transmission $T_{\text{a}}$ as computed in Table 1 can lead to important reduction of the detector’s sensitivity. Such reduction can be partially recovered by tuning other parameters of the coupled filter cavity. However, in a realistic scenario, after the construction of the coupled filter cavity, most of the parameters cannot be adjusted. Thus, deviations in $T_{\text{a}}$ can be partially compensated by tuning only three parameters: the input squeezing angle into the CFC cavities, and the two detunings $\Delta\omega_{a}$ and $\Delta\omega_{c}$ of the sub-cavities [28].
Fig. 3 compares quantum noise reduction for 2FC with the one in the CFC case with a deviation in $T_{\text{a}}$ of 10% and 20% from its nominal value, after partial compensation. We considered all loss sources of Table 1 and we have taken the total cavity length to be $10$ km.
Therefore, by going from 20% deviation, which is the currently achievable tolerance on such a low absolute value of $T_{a}$ [29], to 10%, we can gain up to 2 dB of quantum enhancement above 20 Hz. This underlines the importance of reaching more stringent tolerances on $T_{\text{a}}$. Note that the two planar surfaces (Anti-Reflective and Highly-Reflective) of the middle mirror can be made parallel to each other to form an etalon [30]. Then, the transmissivity can be fine-tuned by controlling the temperature of the mirror, with a precision [31] that should help mitigate imperfections in the initial value of $T_{a}$.
@@
Figure 3: Squeezing degradation for the CFC caused by a 10% and 20% deviation of the middle mirror transmission $T_{\text{a}}$ compared to the CFC model with the optimal $T_{\text{a}}$ and the two FC case. The 10% and 20% $T_{\text{a}}$ deviations are being partly compensated by adjusting the detunings of the CFC sub cavities and the injected squeezing angle. Top: Noise spectral density comparison. Bottom: Quantum enhancement in dB. Losses are being considered as detailed in Table 1.
@@ III.1 Optical Losses
Figure 4: Squeezing degradation for Round Trip Loss $\Lambda$ for 2FC (solid lines) and CFC (dashed lines), for a total length $L_{1}+L_{2}=L_{a}+L_{c}=5\text{km}+5\text{km}=10$ km. From top to bottom: Quantum noise reduction (for 10 dB of input squeezing); Relative quantum noise reduction between CFC and 2FC (a positive value means that CFC performs better than 2FC); the squeezing efficiency $\eta$, the dephasing $\sqrt{\Xi}$ and the misphasing $\Delta\theta_{D}$.
Fig. 4 compares the squeezing degradation when Round Trip Losses $\Lambda$ are added in the CFC and 2FC models (which can be physically interpreted as absorption loss on each mirror), for values of $\Lambda$ equal for all mirrors and ranging from 0 to 30 ppm.

### v2

@@ Abstract
Gravitational-wave detectors use state-of-the-art quantum technologies to reduce the noise induced by vacuum fluctuations, via injection of squeezed states of light. Future detectors, such as Einstein Telescope, may require the use of two filter cavities or a 3-mirror coupled filter cavity to achieve a complex rotation of the squeezing ellipse, in order to reduce the quantum noise over the whole detector bandwidth. In this work, we compare the theoretical feasibility and performances of these two optical layouts and their resilience with respect to different degradation sources (optical losses, mismatching, locking precision), analytically and numerically. We extend previous analysis on squeezing degradation and find that the coupled cavity scheme provides similar or better performances than the two-cavity option, in terms of resilience with respect to imperfections and optical losses. We further highlight the role of mode-mismatch phases in limiting squeezing. Finally, we propose a possible two-step implementation scheme for Einstein Telescope using a single filter cavity that can be possibly upgraded into a coupled filter cavity.
@@ I Introduction
Over the past decade, LIGO, Virgo and KAGRA have opened the era of gravitational-wave (GW) astronomy, detecting hundreds of signals from merging compact binary objects, including black-hole binaries, neutron-star binaries and black-hole-neutron-star systems [1, 2]. These observations have led to a wide range of scientific insights in fields as diverse as general relativity, astrophysics and cosmology [3].
@@
Quantum noise limits the ultimate sensitivity of GW detectors, due to vacuum fluctuations of the electromagnetic field entering the detector through the output port. It can be mitigated by injecting squeezed vacuum through this port [4, 5]. Because of the inherent frequency response of the interferometer, and the ponderomotive coupling with the mirrors, such a squeezed vacuum state needs to acquire a frequency dependence in order to optimally reduce quantum noise in the whole detector bandwidth. Precisely, the interferometer rotates the squeezed quadratures in a frequency-dependent manner [6]. Fig. 1 shows the simplified optical scheme of such a quantum-enhanced GW interferometer.
@@
Figure 1: Simplified optical scheme for dual-recycled GW interferometer with Frequency-Dependent Squeezing injected through the output port via Faraday cirulator/isolator. Current and future generations of GW detectors use the same simplified scheme, but ET-LF operates in a detuned configuration. Here, for clarity, we represent the arms at a right angle.
@@
Frequency-Dependent Squeezing (FDS) is achieved by first generating a frequency-independent squeezed vacuum state, then passing it through one or several filter cavities to rotate its quadratures, thus compensating the response of the interferometer [7, 8, 9, 10]. Recently, the Advanced LIGO detectors achieved FDS using a single filter cavity [11], which improved the detection rate up to 65% [12] and allowed continuous operation below the standard quantum limit [13].
@@
At present, “third-generation” (3G) detectors, to be built in new infrastructures, are being studied and aim to increase sensitivity by a factor of $\sim 10$ compared to existing detectors. They will continue the scientific program of LIGO-Virgo-KAGRA, addressing fundamental open questions such as the nature of gravity and dark energy, the properties of nuclear matter, and the formation of neutron stars and black holes throughout cosmic history [14]. Einstein Telescope, one of such 3G detectors, is composed of two interferometer designs: a “room temperature” detector, optimized for high-frequency sensitivity, and a “cold” detector (ET-LF), operating at cryogenic temperatures for improved low-frequency sensitivity [15]. This work focuses on the quantum noise reduction (QNR) in ET-LF for each 10-km, detuned, dual-recycled Fabry-Pérot Michelson interferometer, keeping the results as general as possible and independent on the exact angle between the arms. We shall only particularize to the triangular geometry currently proposed [15] for Einstein Telescope when computing the astrophysical reach.
ET-LF is expected to operate in a detuned (non-broadband) configuration [16, 17]. This so-called “detuning” corresponds to a constant, $\mu m$-scale offset of the position of the signal recycling mirror (see Fig. 1) away from the resonant length for the carrier field. Physically, this induces a resonant extraction of the GW signal around a frequency determined by the detuning, thus tailoring the detector to be more sensitive to frequency ranges of interest. Due to this complex response of the interferometer, the corresponding optimal FDS scheme requires strictly more than a single filter cavity; usually two (referenced as 2FC in the following [18]) are needed.
@@
Among alternatives schemes that do not require two physical filter cavities, such as EPR squeezing [19, 20, 21] and quantum teleportation [22], we consider here a three-mirror cavity, otherwise known as a coupled filter cavity (henceforth referenced as CFC), which has been first proposed in [23]. A comparison of 2FC vs. CFC in the lossless case was performed in [24], where it was identified that the equivalence between the two schemes imposed a value for the CFC middle mirror transmissivity apparently smaller than current state of the art techniques in mirror coating.
@@
In this paper, we review the equivalency between 2FC and CFC, and address the previous issue of the middle mirror transmission (Section II). In particular, we show that, for longer filter cavities than that studied in [24], its value is actually readily achievable; we find that transmission errors can be compensated by adjusting the cavities’ detunings. We then perform a complete comparison between 2FC and CFC configuration in terms of squeezing degradation budget, in particular including the effects of optical loss, mismatch and phase noise (Section III). It is shown that, theoretically, CFC may provide better resilience to mode-matching and phase noise than 2FC. Finally, we introduce a tuned configuration of ET-LF (with single filter cavity) as an intermediary step before achieving detuned ET-LF, and study its astrophysical merit (Section IV).
@@ II.1 Scaling law between 2FC and CFC
with $\nu_{c}=\frac{c}{2L_{c}}$ the free spectral range (FSR) of the c-cavity, $\gamma_{i}=\frac{cT_{i}}{4L_{i}}$ and $\Delta\omega_{i}$ the linewidth and detuning of cavity labeled $i$, where $T_{i}$ is the input mirror transmissivity of the $i$-th cavity.
@@
Table 1 provides the parameters which will be used in the rest of this paper, generally similar to previous studies [24]. We briefly comment on the chosen value of 30 ppm for round trip loss per cavity $\Lambda$. They have been estimated according to the empirical scaling law:
$\Lambda=30\text{ ppm}\times\left(\frac{L}{300\text{ m}}\right)^{0.3}\left(\frac{1064\text{ nm}}{\lambda}\right)^{2}$ (5)
where the values of 30 ppm has been found by simulations in [25] for a 300-m cavity illuminated with a 1064 nm laser. For the scaling law with the length of the cavity we use the result of [26]. The scaling with the laser wavelength is given by the physics of the scattering [27]. This can be considered a conservative estimation, since we expect future advances in the mirror and coating techniques.
@@
$\Lambda_{\text{in}}$ Injection Faraday loss $0.01$
$\Lambda_{\text{arm}}$ Arm round-trip loss $45$ ppm
$\Lambda_{\text{SR}}$ SEC round trip loss $1000$ ppm
$\theta_{\text{HD}}$ Homodyne angle 0.27-0.27 rad
@@
$\Lambda$ Round Trip Loss per FC 30 ppm
$\Lambda_{\text{sub}}$ Substrate Loss (CFC) 20 ppm
$\delta L$ FC RMS length noise 1 pm
@@
$L_{\text{FC1,2}}$ Length of each FC 55 km
@@
$L_{\text{FC1,2}}$ Length of each FC 11 km
$L_{\text{CFC}}$ Total length of CFC 22 km
@@
Table 1: Parameters used in this study. For comparison, we show the parameters of 2FC and CFC if the total length is either 2 or 10 km. The round trip losses $\Lambda_{\text{arm}},\Lambda_{\text{SR}},\Lambda$ are computed per cavity of the interferometer (Arm or Signal Extraction Cavity) and frequency-dependent systems (2FC or CFC); the substrate loss $\Lambda_{\text{sub}}$ for the CFC case is accounted separately.
@@ II.2 The middle mirror in CFC
The middle mirror is a critical component of the CFC configuration. We first comment on concerns raised by the unusually small value for $T_{a}$. First, note that while its value in the $L_{\text{CFC}}=2$ km case ($T_{a}<0.3$ ppm) is indeed much lower than what is implemented in the current generation of GW detectors, the $L_{\text{CFC}}=10$ km case ($T_{a}=6.75$ ppm) is comparable to the transmission of the end mirror of the currently installed filter cavity in Virgo (see [28]), thus technically already achieved for wavelengths in the infrared spectrum. Therefore, in the rest of this paper, unless explicitly specified otherwise, we will work with $L_{\text{CFC}}=10$ km, which seemingly is within technological reach.
Interestingly, if one considers the a-cavity as a single cavity – that is, the input mirror is $T_{a}$ – then it may seem that such a cavity is loss-dominated ($\Lambda=30\text{ ppm}\gg T_{a}$), and that the finesse would be of the order of $10^{5}$. However, since the a-cavity is embedded in a coupled cavity system, the correct qualitative picture is to consider that the first two mirrors ($T_{c}$ and $T_{a}$) form a cavity, which is equivalent to an effective mirror with variable transmission $T_{\text{eff}}$ given as a function of the sideband frequency $\Omega$ by
@@
with $r=\sqrt{R_{a}R_{c}(1-\Lambda)}$. Since $\Delta\omega_{c}=12$ Hz and we are interested in sideband frequencies in the range of the expected GW signals $\Omega\in 2\pi\times[-30\text{ Hz},30\text{ Hz}]$, we have that $\Omega-\Delta\omega_{c}\in 2\pi\times[-42\text{ Hz},18\text{ Hz}]$. Plugging this into the previous formula, we find that $T_{\text{eff}}\in[210\text{ ppm},\ 10600\text{ ppm}]$. This means that the effective second cavity is never loss-dominated in the range of frequencies of interest. The associated finesse range is $\mathcal{F}\in[6\cdot 10^{2},\ 3\cdot 10^{4}]$ , whose upper value’s order of magnitude is comparable to that of current squeezing filter cavities in LIGO & Virgo ($\mathcal{F}\sim 10^{4}$, see [11, 7]).
Technical limitations in the design of low-transmissivity mirrors may lead to small deviations from the theoretical optimal value $T_{\text{a}}$, as computed in Table 1, which in turn imply important reductions of the GW detector’s sensitivity. These deviations can be directly or indirectly compensated, respectively through thermal actuators on the mirror or detuning offsets on the cavities’ resonant frequencies.
Thermal controls enable direct tunability of $T_{a}$ in the following way: the two planar surfaces (Anti-Reflective and Highly-Reflective) of the middle mirror can be made parallel to each other to form an etalon [29]. Then, a controlled variation in the temperature of the mirror substrate changes the substrate index of refraction (thermo-refractive effect) as well as the physical length of the substrate (thermo-expansion), thus modifying the optical path length inside of the mirror [30]. This can usually be done with a precision that should help mitigate imperfections in the initial value of $T_{a}$. Denoting $R_{\text{AR}}$ the reflectivity of the AR surface, we find that the transmission as a function of substrate temperature offset $\Delta\theta$ from room temperature has the form $T_{a}(\Delta\theta)\simeq T_{a}[1+2\sqrt{R_{\text{AR}}}\cos(2\pi\Delta\theta/\Delta\theta_{T})]$, where $\Delta\theta_{T}=1.49$ K for typical values in Fused Silica substrates. This means that a choice of $R_{\text{AR}}=0.01$ allows to compensate a $20\%$ deviation of the value $T_{a}$ by tuning the mirror temperate. With a typical temperature precision of $\Delta\theta_{\text{min}}\sim 10$ mK, this scheme can control $T_{a}$ down to a relative error of $|\Delta T_{a}/T_{a}|\sim 1\%$, which is more than an order of magnitude improvement and ensures that the residual degradation of the GW detector sensitivity is negligible.
If thermal controls are unavailable, it is still possible to partially recover the detector sensitivity by tuning other parameters of the coupled filter cavity. However, in a realistic scenario, after the construction of the coupled filter cavity, most of the parameters cannot be adjusted. Thus, deviations in $T_{\text{a}}$ can be partially compensated by tuning only three parameters: the input squeezing angle into the CFC cavities, and the two detunings $\Delta\omega_{a}$ and $\Delta\omega_{c}$ of the sub-cavities [31]. In practice, these detunings are controlled by locking the cavity on resonance with an auxiliary laser, which is frequency-locked at an offset from the main laser. By precisely controlling this offset, one can arbitrarily adjust the resonance conditions of the cavity and thus its frequency detuning from the squeezed field. Fig. 3 compares quantum noise reduction for 2FC with the one in the CFC case with a deviation in $T_{\text{a}}$ of 10% and 20% from its nominal value, after such partial compensation. We considered all loss sources of Table 1 and we have taken the total cavity length to be $10$ km. Therefore, by going from 20% deviation (dashed red), which is the currently achievable tolerance on such a low absolute value of $T_{a}$ [32], to 10% (dashed green), we can gain up to 1.5 dB of quantum enhancement above 30 Hz. This underlines the importance of reaching stringent tolerances on $T_{\text{a}}$.
@@
Figure 3: Squeezing degradation for the CFC caused by 10% (dashed green) or 20% (dashed red) deviations of the middle mirror’s transmission $T_{\text{a}}$, compared to the CFC model with the optimal $T_{\text{a}}$ (solid red) and the two FC case (blue). The sensitivity without squeezing (solid black) is plotted for reference. The 10% and 20% $T_{\text{a}}$ deviations are being partly compensated by adjusting the detunings of the CFC sub-cavities and the injected squeezing angle. Losses are detailed in Table 1 and Section III.4. Top: Noise spectral density comparison for a single L-shaped interferometer. Bottom: Quantum enhancement in dB.
@@ III.1 Optical Losses
Figure 4: Squeezing degradation for Round Trip Loss per cavity $\Lambda$ for 2FC (solid lines) and CFC (dashed lines), for a total length $L_{1}+L_{2}=L_{a}+L_{c}=5\text{km}+5\text{km}=10$ km. From top to bottom: Quantum noise reduction (for 10 dB of input squeezing); Relative quantum noise reduction between CFC and 2FC (a positive value means that CFC performs better than 2FC); the squeezing efficiency $\eta$, the dephasing $\sqrt{\Xi}$ and the misphasing $\Delta\theta_{D}$.
Fig. 4 compares the squeezing degradation when Round Trip Losses $\Lambda$ are added in the CFC and 2FC models (which can be physically interpreted as absorption loss on each mirror), for values of $\Lambda$ equal for all mirrors and ranging from 0 to 30 ppm per cavity. In the case of CFC, this excludes additional substrate loss due to passing through the middle mirror to go from one sub-cavity to the other. The substrate loss are considered separately in the full degradation budget.